I need help on these two questions!
1. Write the standard form of the equation of the line passing through the point (–5, 4) and perpendicular to the line - 3x - 2y = 8.
2. Find the slope-intercept form of the the line passing through the point (1,3) and parallel to the line y = 5x + 3
THANKS!

Answer by sachi(548)

1. Write the standard form of the equation of the line passing through the point (–5, 4) and perpendicular to the line - 3x - 2y = 8.
2. Find the slope-intercept form of the the line passing through the point (1,3) and parallel to the line y = 5x + 3
ans 1:
the line is perpendicular to - 3x - 2y = 8 or y=-3/2x-4 with slope =-3/2
so the required slope of the line=2/3
& passes through (–5, 4)
so the eqn is [y-4]/[x-(-5)]=2/3
or 3y-12=2x+10
or 3y=2x+22 ans
ans 2:
the line is parallel to the line y = 5x + 3 with slope =5
so the required slope of the line=5
& passes through (1,3)
so the eqn is [y-3]/[x-1]=5
or y-3=5x-5
or y=5x-2 ans